# Alexandrov spaces are CS sets

### v1

@@ 2.5. Extremal subsets
Let $\Sigma$ be the Poincaré homology $3$-sphere of constant curvature $1$. Then the spherical suspension $S(\Sigma)$ is not a topological manifold, but the double spherical suspension $S^{2}(\Sigma)$ is a topological sphere by the theorem of Cannon [7] and Edwards [9], [10]. Thus the MCS stratification of $S^{2}(\Sigma)$ is trivial. On the other hand, since $\operatorname{diam}\Sigma\leq\pi/2$ (cf. [16]), the extremal stratification consists of the subsuspension $S(\{\xi,\eta\})$ and its complement, where $\xi,\eta$ are the poles of $S(\Sigma)$.
@@
Theorem 1.2 tells us that the above proper extremal subset must exist. See Theorem 4.1 and Proposition 4.4 for further discussions.
@@
Now we prove Theorem 1.1. Indeed, almost all the background knowledge needed for the proof is covered in the previous section.
@@
Let $X$ be an Alexandrov space and $E$ a primitive extremal subset. We will show that there exists a compact stratified set $L$ (depending only on $E$) such that for any point $x$ in the main part $\mathring{E}$ of $E$ there is a neighborhood $U$ of $x$ in $X$ admitting an isomorphism as stratified sets
$U\cong B^{m}\times C(L),$
where $m=\dim E$ and the stratifications of the left- and right-hand sides are the natural ones induced by those of $X$ and $L$, respectively.
Step 1. We first prove the claim for any regular point of $E$. A regular point of $E$ is a point where the tangent cone of $E$ is isometric to Euclidean space of dimension $m=\dim E$ (the tangent cone means the Gromov–Haussdorff blowup of $E$ at $p$). Note that regular points are contained in $\mathring{E}$. See [12] for more information; but actually all we need here is the second half of Remark 2.16.
Let $p\in E$ be a regular point. Then there exists a regular map $f$ to $\mathbb{R}^{m}$ at $p$ (consider the distance map from $m$ points in $E$ close to $p$ that make obtuse but almost right angles at $p$). Since $\dim E=m$, there is no function $f_{1}$ such that $(f,f_{1})$ is regular at $p$ (Remark 2.16). By Theorem 2.8 and Remark 2.10, one can construct a function $g$ and the associated canonical neighborhood $K=K(f,g)$ of $p$ with a homeomorphism
$h:K\to B^{m}\times C(L_{p}),$
@@
It remains to show that $h$ preserves the stratifications of both sides. Note that not only $K$, but also $L_{p}$ has a natural stratification as a subset of $X$, by restricting the extremal stratification. Since $L_{p}$ is not open, we have to show that this is a stratified set in the sense of Section 2.1, i.e., the components of the strata are open. However, we defer it to the next step and first observe that $h$ preserves these stratifications.
By Complement 2.19, $h$ can be chosen to preserve extremal subsets. Thus the only thing we should check here is that there is a bottom stratum in the left-hand side corresponding to $B^{m}\times\{o\}$ in the right-hand side. This is nothing but $E$ itself by the same reasoning $\dim E=m$ as above (Theorem 2.8(2), Remarks 2.16).
Step 2. As mentioned above, we have to prove that the components of the strata of $L_{p}$ are open. It is sufficient to show that the strata are locally path-connected. Let $E_{1}$ be a primitive extremal subset intersecting $L_{p}$ (thus containing $E$) and let $x\in L_{p}\cap\mathring{E}_{1}$. By Theorem 2.8 and Remark 2.10, there exists a canonical neighborhood $K_{1}$ of $x$ respecting $(f,g)$ (see Definition 2.11). By Complement 2.19, it follows that $K_{1}\cap L_{p}\cap\mathring{E}_{1}$ is homeomorphic to a product of a ball and a cone, which is locally path-connected at $x$.
Step 3. Finally we prove the claim for any point of the main part $\mathring{E}$ with fixed $L$. Fix a regular point $p\in E$ and set $L:=L_{p}$. Define a subset $F$ of $E$ by
@@
where the isomorphism is as stratified sets. Note that $F$ is closed since the bottom stratum of $B^{m}\times C(L)$ has dimension $m$, while $\mathring{E}$ is the unique $m$-dimensional stratum contained in $E$. However, we will not need it in view of Lemma 2.17 used below (but see also Remark 3.1, where it is needed).
We prove that $F$ is an extremal subset of $X$. It suffices to verify the condition (*) of Lemma 2.17 for $F$. Suppose $x\in E$ has a (small) neighborhood isomorphic to $B^{m}\times C(L)$. For a canonical neighborhood $K(f_{1},g_{1})$, we show that if $x\in g_{1}^{-1}(0)$, then any other $y\in g_{1}^{-1}(0)$ is contained in $E$ and has a neighborhood isomorphic to $B^{m}\times C(L)$. This follows from Complement 2.19 (clearly $y\in E$ by Remark 2.18). Indeed, since $x,y\in g_{1}^{-1}(0)$, there is a homeomorphism between neighborhoods of $x$ and $y$ via a translation with respect to the $f_{1}$-coordinate of $K(f_{1},g_{1})$. Since the homeomorphism of Complement 2.19 respects extremal subsets, so does the translation. This implies that $x$ and $y$ have isomorphic neighborhoods. Thus the condition (*) holds and hence $F$ is an extremal subset of $X$.
Moreover, $F$ is not equal to $E$ since $p\notin F$. By the primitiveness, $F$ must be contained in $E\setminus\mathring{E}$. Therefore, any point of $\mathring{E}$ has a neighborhood isomorphic to $B^{m}\times C(L)$. This complete the proof. ∎
@@
Actually in Step 3, we do not need Lemma 2.17 to prove that $F$ is an extremal subset. One can just repeat the same argument as in [29, 3.8, 1.2], which showed that $\mathring{E}$ is a topological manifold, replacing the absolute fibration theorem on extremal subsets in [29, §2] by the relative one in [20, §9]. The reason we did not take this approach is because our exposition puts emphasis on canonical neighborhoods (in fact, we have not even given the precise definition of extremal subsets). Another reason is that our argument can be viewed as a fixed space version of [13, 5.6, 5.11], which showed that the homotopy type of a regular fiber in a collapsing space is invariant over a component of a stratum of the limit space.
By the theorem of Handel [18, 2.4], we obtain the following corollary.
@@
Here the intrinsic stratification (or minimal stratification) is defined as follows. The $k$-dimensional stratum consists of those points whose conical neighborhood topologically splits into $B^{k}\times C(L)$ but not $B^{k+1}\times C(L^{\prime})$, where $L$ and $L^{\prime}$ are compact topological spaces. Compare with the MCS stratification in Section 2.3, where $L$ and $L^{\prime}$ are MCS spaces. Although the MCS stratification might be less important now, the following question still remains.
@@
Another question remains regarding the link, which leads to Theorem 1.2.
@@
Note that $L$ is an MCS space (Theorem 2.8). Moreover, in view of the stability theorem [25, 4.3], $L$ will be homeomorphic to a space $\Sigma$ of curvature $\geq 1$, which is the subspace of directions at $p$ normal to the tangent cone of $E$. Therefore $L$ admits a CS structure by the extremal stratification of $\Sigma$. However, at present it is unclear to the author whether the above stratification of $L$ induced from the extremal stratification of the ambient space $X$ has such nice properties. In view of the relative stability theorem [20, 9.2], this stratification will exist in between the first MCS and the second CS stratifications mentioned above.
@@
Next we prove Theorem 1.2. Here we will assume some more familiarity with Alexandrov geometry. The following are basic properties of Alexandrov topological manifolds and homology manifolds (with integer coefficients). We refer to [35], [23], and [33, 7.2]: the second one is concerned with geodesically complete spaces with curvature bounded above, but the methods there can also be applied to Alexandrov spaces with curvature bounded below (if correctly translated).
  • Any space of directions of an Alexandrov homology manifold is a homology manifold with the homology of a sphere.
  • Any nonmanifold point of an Alexandrov homology manifold is isolated; in particular, it is an extremal point.
  • A point in an Alexandrov homology manifold is a manifold point if and only if its space of directions is simply-connected, or equivalently homotopy equivalent to a sphere (by the Whitehead theorem).
Moreover, in dimension $\leq 4$, any space of directions of an Alexandrov topological manifold is homeomorphic to a sphere. Therefore we may restrict our attention to dimension $\geq 5$. We will prove the following slightly stronger statement.
@@
Let $X$ be an Alexandrov topological manifold. Let $E$ be the set of points in $X$ where the space of directions is not homeomorphic to a sphere. If $E$ is nonempty, then it is a $1$-dimensional extremal subset.
Since an Alexandrov space without proper extremal subsets is a topological manifold, this immediately implies Theorem 1.2.
@@
We denote by $\Sigma_{p}$ the space of directions at $p\in X$. Since $X$ is a topological manifold, $\Sigma_{p}$ is a homology manifold with the homotopy type of a sphere as stated above. In particular, $\Sigma_{p}$ is homeomorphic to a sphere if and only if it is a topological manifold (by the generalized Poincaré conjecture). Together with the above-mentioned properties, we see that
  • $\Sigma_{p}$ is a topological sphere iff its all spaces of directions are simply-connected;
  • the nonmanifold points of $\Sigma_{p}$, or equivalently, points having nonsimply-connected spaces of directions, are isolated extremal points.
These facts will be used implicitly and frequently below.
If $p\in E$, we also denote by $\Sigma_{p}E$ the space of directions of $E$ at $p$, that is, the subset of $\Sigma_{p}$ consisting of limits of the directions of shortest paths from $p$ to $p_{i}\in E\setminus\{p\}$ that converges to $p$. To prove that $E$ is an extremal subset, it is sufficient to show that $E$ is closed and $\Sigma_{p}E$ is an extremal subset of $\Sigma_{p}$ for any $p\in E$ ([29, 1.4]). For the latter, we will prove that $\Sigma_{p}E$ coincides with the set of nonmanifold points in $\Sigma_{p}$, which is a discrete set of extremal points as stated above (but see also Step 1 below).
Step 1. We first show that if $\Sigma_{p}$ contains only one nonmanifold point $\xi$, then $\Sigma_{p}$ is contained in the closed $\pi/2$-neighborhood of $\xi$. This is an additional condition imposed on a one-point extremal subset in the space of directions so that the above infinitesimal characterization [29, 1.4] holds (see [29, 1.1], [31, §4.1 Property (2)]). Actually we prove that $\operatorname{diam}\Sigma_{p}\leq\pi/2$.
Suppose $\operatorname{diam}\Sigma_{p}>\pi/2$. Then the suspension theorem of Perelman [25, 4.5] implies that $\Sigma_{p}$ is homeomorphic to a suspension over a space of curvature $\geq 1$. Moreover, its proof shows that the distance function to each pole of the suspension is regular except at the poles. In particular, any point other than the poles is not an extremal point, thus a manifold point. Therefore $\xi$ must be one of the poles of the suspension, and the other pole, which also has a nonsimply-connected link, is a nonmanifold point. This is a contradiction.
Step 2. We next prove the following
@@
Let $p\in X$ and $\xi\in\Sigma_{p}$. Then $\xi$ is a nonmanifold point if and only if there exists a sequence $p_{i}\in E\setminus\{p\}$ converging to $p$ such that the directions to $p_{i}$ converges to $\xi$.
@@
We denote by $B(p,r)$ and $S(p,r)$ the open $r$-ball and $r$-sphere around $p$, respectively. First suppose $r$ is small enough and let $x\in S(p,r)$. Then the point $p$ is a $1$-strainer at $x$, that is, there exists a direction at $x$ making an angle almost $\pi$ with the directions to $p$, which we call an almost antipode. Next choose $\varepsilon>0$ small enough compared to $r$. Then the pair of points $(p,x)$ is a $2$-strainer at any point of $S(p,r)\cap B(x,\varepsilon)\setminus\{x\}$, i.e., the directions to $p$ and $x$ and their almost antipodes are almost orthogonal. By the stability theorem and the suspension theorem [25, 4.3, 4.5], we see that $\Sigma_{x}$ is homeomorphic to the suspension over $\Sigma:=S(p,r)\cap S(x,\varepsilon)$ (more precisely, the direction(s) at $x$ to $p$ may not be a pole of the suspension, but they are very close to the pole, so the homeomorphism type of the equator of the suspension turns out to be $\Sigma$). Moreover, as mentioned in Step 1, any point other than the poles of the suspension is not an extremal point, thus a manifold point. Therefore, $\Sigma_{x}$ is a topological manifold if and only if $\Sigma$ is simply-connected.
@@
Since $r$ is small enough, it follows from the stability theorem that $S(p,r)$ is homeomorphic to $\Sigma_{p}$; in particular, it is a homology manifold (of dimension $\geq 4$) with isolated nonmanifold points. Since $\varepsilon$ is small enough, it also follows from the fibration theorem (with a gluing argument) that $S(p,r)\cap B(x,\varepsilon)$ is homeomorphic to the cone over $\Sigma$. Thus $\Sigma$ is simply-connected if and only if $x$ is a manifold point of $S(p,r)$ (the “only if” part follows from [8], [5]; see also [23, 6.2]). Since the rescaled sphere $r^{-1}S(p,r)$ converges to $\Sigma_{p}$ as $r\to 0$, and the stability homeomorphism is a Gromov–Hausdorff approximation, the claim follows. ∎
Now the “if” part of Claim 4.2 implies that $E$ is closed and $\Sigma_{p}E$ is contained in the set of nonmanifold points of $\Sigma_{p}$. On the other hand, the “only if” part implies that the nonmanifold points of $\Sigma_{p}$ are contained in $\Sigma_{p}E$. This complete the proof of Theorem 4.1. ∎

### v2

@@ 2.6. Double suspension
Let $\Sigma$ be the Poincaré homology sphere of constant curvature $1$. Then we have
  • $S(\Sigma)$ is not a topological manifold;
  • $S^{2}(\Sigma)$ is a topological sphere by Cannon [7] and Edwards [9], [10].
In particular, the MCS stratification of $S^{2}(\Sigma)$ is trivial. On the other hand, since $\operatorname{diam}\Sigma\leq\pi/2$ (cf. [16]), the extremal stratification consists of the subsuspension $S(\{\xi,\eta\})$ and its complement, where $\xi,\eta$ are the poles of $S(\Sigma)$. Notice that the extremal subset $S(\{\xi,\eta\})$ coincides with the set of points with nonspherical links.
@@
Theorem 1.2 tells us that the above extremal subset must exist. See Theorem 4.1 and Corollary 4.6 for further discussion.
@@
Now we prove Theorem 1.1. Almost all the background knowledge needed for the proof is covered in the previous section.
@@
Let $X$ be an Alexandrov space and $E$ a primitive extremal subset of dimension $m$. We will show that there exists a compact stratified set $L$ (depending only on $E$) satisfying the following: for any point $x$ in the main part $\mathring{E}$ of $E$, there is a neighborhood $U$ of $x$ in $X$ that admits an isomorphism as stratified sets
$U\cong B^{m}\times C(L).$ (3.1)
Here the stratifications of the left- and right-hand sides are the natural ones induced from those of $X$ and $L$, respectively.
Step 1. We first prove (3.1) for any regular point of $E$. A point $p\in E$ is called regular if the tangent cone $T_{p}E$ of $E$ at $p$ is isometric to $\mathbb{R}^{m}$, where $m=\dim E$. Here the tangent cone means the Gromov–Hausdorff blowup; see [11] for the details. Note that the set of regular points of $E$ is dense in $E$ and is contained in $\mathring{E}$ ([11, 1.1], Remark 2.20).
Let $p\in E$ be a regular point. Choose $(m+1)$ directions $\xi_{i}$ ($0\leq i\leq m$) in the tangent cone $T_{p}E=\mathbb{R}^{m}$ that make obtuse angles at the vertex and let $a_{i}\in X$ be points near $p$ in the directions $\xi_{i}$. Then the distance map
$f=(d(a_{1},\cdot),\dots,d(a_{m},\cdot))$
is regular at $p$. Since $\dim E=m$, there is no function $f_{1}$ such that $(f,f_{1})$ is regular at $p$ (Remark 2.20). By Theorem 2.10 and Remark 2.12, one can construct a function $g$ and a canonical neighborhood $K=K(f,g)$ of $p$ together with a homeomorphism
$h:K\to B^{m}\times C(L_{p}),$ (3.2)
@@
It remains to show that $h$ preserves the stratifications of both sides of (3.2). Note that not only $K$, but also $L_{p}$ has a natural stratification as a subset of $X$, by restricting the extremal stratification of $X$. Since $L_{p}$ is not open, we have to show that this is a stratified set in the sense of Section 2.1, i.e., the components of the strata are open. However, we defer it to the next step and first observe that $h$ preserves these stratifications.
By Complement 2.23, $h$ can be chosen to respect all extremal subsets. Note that the right-hand side of (3.2) may have one extra stratum $B^{m}\times\{o\}$ arising from the cone structure of $C(L_{p})$. We have to show that there is a bottom stratum in the left-hand side that corresponds to $B^{m}\times\{o\}$. This is nothing but $E$ itself by the same reasoning $\dim E=m$ as above (Theorem 2.10(2), Remark 2.20).
Step 2. As mentioned above, we prove that the components of the strata of $L_{p}$ are open. It suffices to show that the strata are locally path-connected. Let $E_{1}$ be a primitive extremal subset of $X$ intersecting $L_{p}$ and let $x\in L_{p}\cap\mathring{E}_{1}$. By Theorem 2.10 and Remark 2.12, there exists a canonical neighborhood $K_{1}$ of $x$ respecting $(f,g)$ (see Definition 2.13). By Complement 2.23, $K_{1}\cap L_{p}\cap\mathring{E}_{1}$ is homeomorphic to a product of a ball and a cone, which is locally path-connected at $x$.
Step 3. Finally we prove (3.1) for any point of the main part $\mathring{E}$. Fix a regular point $p\in E$ and set $L:=L_{p}$. Define a subset $F$ of $E$ by
@@
where the isomorphism is as stratified sets. Note that $F$ is closed since the bottom stratum of $B^{m}\times C(L)$ has dimension $m$, while $\mathring{E}$ is the unique $m$-dimensional stratum contained in $E$.
We prove that $F$ is an extremal subset of $X$. It suffices to verify Condition (*) of Lemma 2.21 for $F$. Suppose $x\in E$ has a small neighborhood isomorphic to $B^{m}\times C(L)$. For a canonical neighborhood $K(f_{1},g_{1})$, we show that if $x\in g_{1}^{-1}(0)$, then any $y\in g_{1}^{-1}(0)$ has a neighborhood isomorphic to $B^{m}\times C(L)$ (note that $y\in E$ follows from Remark 2.22). Consider a homeomorphism
$h_{1}:K(f_{1},g_{1})\to B^{k}\times C(L_{1})$
respecting extremal subsets as in Complement 2.23 and a homeomorphism
$T_{1}:B^{k}\times C(L_{1})\to B^{k}\times C(L_{1})$
that sends $h_{1}(x)=(f_{1}(x),o)$ to $h_{1}(y)=(f_{1}(y),o)$ and fixes the $C(L_{1})$-coordinate. The composition $h_{1}^{-1}\circ T_{1}\circ h_{1}$ shows that $x$ and $y$ have isomorphic neighborhoods. Thus Condition (*) is satisfied and hence $F$ is an extremal subset of $X$.
Since $p\notin F$, $F$ is a proper subset of $E$. By the definition of the main part, $F$ is contained in $E\setminus\mathring{E}$. Therefore, any point of $\mathring{E}$ has a neighborhood isomorphic to $B^{m}\times C(L)$. This complete the proof. ∎
@@
In Step 3, we do not really need Lemma 2.21 to prove that $F$ is an extremal subset. One can just repeat the same argument as in [33, 3.8, 1.2], which showed that $\mathring{E}$ is a topological manifold, by replacing the absolute fibration theorem on extremal subsets [33, §2] with the relative one [23, §9]. The reason we did not take this approach is because our exposition puts emphasis on canonical neighborhoods. Our argument can be viewed as a fixed space version of [12, 5.6, 5.11], which showed that the homotopy type of a regular fiber in a collapsing Alexandrov space is invariant over a component of the extremal strata of the limit space.
By the theorem of Handel [19, 2.4], we obtain the following corollary.
@@
Here the intrinsic stratification (or minimal stratification) is defined as follows. The $k$-dimensional stratum consists of those points whose conical neighborhood topologically splits into $B^{k}\times C(L)$ but not $B^{k+1}\times C(L^{\prime})$, where $L$ and $L^{\prime}$ are compact topological spaces. Compare with the MCS stratification in Section 2.3, where $L$ and $L^{\prime}$ are MCS spaces. The following question still remains.
@@
Another question remains regarding the link (cf. Theorem 1.2, Conjecture 1.3).
@@
Note that $L$ is an MCS space (Theorem 2.10). Moreover, in view of the stability theorem [29, 4.3], $L$ is homeomorphic to the set $\Sigma$ of directions at $p$ normal to the tangent cone $T_{p}E$. Since $p$ is a regular point, $\Sigma$ has curvature $\geq 1$. Thus $\Sigma$ admits the extremal stratification, which gives $L$ a CS structure. However, this CS stratification is generally finer than the above stratification of $L$ induced from the extremal stratification of the ambient space $X$. See Example 5.7.
@@
Next we prove Theorem 1.2. Here we assume more familiarity with Alexandrov geometry and the properties of extremal subsets.
We first collect the fundamental properties of Alexandrov homology manifolds (with integer coefficients) that will be used throughout this section. By a homology $n$-manifold we mean a topological space whose local homology group is isomorphic to that of $\mathbb{R}^{n}$.
Let $X$ be an Alexandrov homology manifold and $p\in X$.
  1. (1)
    The space of directions $\Sigma_{p}$ at $p$ is an Alexandrov homology manifold with the homology of a sphere.
  2. (2)
    The set of nonmanifold points in $X$ is discrete; in particular, if $p$ is a nonmanifold point, then it is an extremal point.
  3. (3)
    The point $p$ is a manifold point if and only if $\Sigma_{p}$ is simply-connected, or equivalently, homotopy equivalent to a sphere (by (1) and the Whitehead theorem).
  4. (4)
    Every Alexandrov homology manifold of dimension $\leq 3$ is a topological manifold.
For the proofs, we refer to [40] (cf. [41]). More precisely, Properties (1)–(4) can be found in [40, 3.1, 3.5, 1.2, 3.3], respectively. Most of these properties also follow from Theorem 1.1 and [21], [37] (the author thanks Mohammad Alattar for this information). See also [38, 7.2] and the subsequent paragraph. We remark that the corresponding results for geodesically complete spaces with curvature bounded above can be found in [26]. The proofs there, which do not rely on strong structural results as in the Alexandrov case, also apply to Alexandrov spaces with curvature bounded below (if correctly translated).
We now prove Theorem 1.2. We show the following more rigid result.
@@
Let $X$ be an Alexandrov topological manifold. Let $E$ be the set of points in $X$ at which the space of directions is not homeomorphic to a sphere. If $E$ is nonempty, then it is a $1$-dimensional extremal subset.
Since any Alexandrov space without proper extremal subsets is a topological manifold, Theorem 4.1 immediately implies Theorem 1.2.
In order to prove Theorem 4.1, we recall the infinitesimal characterization of extremal subsets proved in [33, 1.4] (cf. [35, §4.1 Property (2)]). Let $X$ be an Alexandrov space and $E$ a subset of $X$ (it need not be extremal or even closed). For any point $p$ in the closure of $E$, we denote by $\Sigma_{p}E$ the space of directions of $E$ at $p$, that is, the subset of $\Sigma_{p}$ consisting of limits of the directions of shortest paths from $p$ to $p_{i}\in E\setminus\{p\}$, where $p_{i}$ converges to $p$. Note that $\Sigma_{p}E$ is always closed, while $E$ may be not.
Lemma 4.2 ([33, 1.4]).
Let $X$ be an Alexandrov space and $E$ a closed subset. Then $E$ is extremal in $X$ if and only if $\Sigma_{p}E$ is extremal in $\Sigma_{p}$ for any $p\in E$ and satisfies the following additional conditions:
  1. (1)
    if $\Sigma_{p}E=\emptyset$, then we have $\operatorname{diam}\Sigma_{p}\leq\pi/2$;
  2. (2)
    if $\Sigma_{p}E=\{\xi\}$, then we have $\Sigma_{p}=\bar{B}(\xi,\pi/2)$.
Here $\operatorname{diam}$ denotes the diameter and $\bar{B}(\cdot,\cdot)$ denotes a closed metric ball.
@@
By Properties (1), (3) and (4), in dimension $\leq 4$, every space of directions of an Alexandrov topological manifold is automatically homeomorphic to a sphere (via the Poincaré conjecture). Therefore we may restrict our attention to dimension $\geq 5$.
Let $X$ be an Alexandrov topological manifold and $p\in X$. By Properties (1) and (3), the space of directions $\Sigma_{p}$ is a homology manifold with the homotopy type of a sphere. In particular, $\Sigma_{p}$ is homeomorphic to a sphere if and only if it is a topological manifold (by the generalized Poincaré conjecture). By Properties (1) and (3) again, $\Sigma_{p}$ is a topological manifold if and only if all its space of directions are simply-connected. In summary,
$\Sigma_{p}$ is a topological sphere
$\iff$ $\Sigma_{p}$ is a topological manifold
[…]
Furthermore, if $\Sigma_{p}$ has nonmanifold points, then they are isolated extremal points (Property (2)). These facts will be used frequently below.
To prove that $E$ is an extremal subset, it is sufficient to show that
  • $E$ is closed;
  • $\Sigma_{p}E$ is an extremal subset of $\Sigma_{p}$ for any $p\in E$;
  • $\Sigma_{p}E$ satisfies the additional conditions as in Lemma 4.2.
In particular, for the second claim, we will prove in Claim 4.3 that $\Sigma_{p}E$ coincides with the set of nonmanifold points in $\Sigma_{p}$, which is a discrete set of extremal points (Property (2)).
Step 1. We first show that if $\Sigma_{p}$ contains only one nonmanifold point $\xi$, then $\Sigma_{p}$ is contained in the closed $\pi/2$-neighborhood of $\xi$. This is an additional condition required to apply Lemma 4.2 (as we will see below, we do not have to consider the other case $\Sigma_{p}E=\emptyset$). Actually we prove that $\operatorname{diam}\Sigma_{p}\leq\pi/2$.
Suppose $\operatorname{diam}\Sigma_{p}>\pi/2$. Then the suspension theorem of Perelman [29, 4.5] implies that $\Sigma_{p}$ is homeomorphic to a suspension $S(\Sigma)$, where $\Sigma$ is a space of curvature $\geq 1$. Moreover, its proof shows that the distance function to each pole of $S(\Sigma)$ is regular except at the poles. In particular, any point other than the poles is not an extremal point, and thus a manifold point (Property (2)). Therefore the nonmanifold point $\xi$ must be one of the poles of $S(\Sigma)$. Let $\eta$ be the other pole. By the suspension structure and the stability theorem [29, 4.3], we see that $\Sigma_{\eta}$ is homotopy equivalent to $\Sigma_{\xi}$ (more precisely, they are both homotopy equivalent to $\Sigma$). Since $\xi$ is a nonmanifold point, this implies that $\eta$ is also a nonmanifold point (Property (3)). This contradicts the assumption that $\xi$ is the only nonmanifold point of $\Sigma_{p}$.
Step 2. We next prove the following key result.
@@
Let $p\in X$ and $\xi\in\Sigma_{p}$. Then $\xi$ is a nonmanifold point if and only if there exists a sequence $p_{i}\in E\setminus\{p\}$ converging to $p$ such that the directions of shortest paths from $p$ to $p_{i}$ converges to $\xi$. In other words,
$\Sigma_{p}E=N(\Sigma_{p}),$
where $N(\Sigma_{p})$ denotes the set of nonmanifold points in $\Sigma_{p}$.
The idea of the proof is as follows. Suppose a neighborhood of $p$ is isometric to a metric cone over $\Sigma_{p}$. Let $x\in X$ be a point in the direction $\xi\in\Sigma_{p}$. Then $\Sigma_{x}$ is isometric to the spherical suspension $S(\Sigma_{\xi})$, where $\Sigma_{\xi}$ is the space of directions of $\Sigma_{p}$ at $\xi$. As in Step 1, $\Sigma_{x}$ is homeomorphic to a sphere if and only if $\Sigma_{\xi}$ is simply-connected. The former is equivalent to $\xi\in\Sigma_{p}E$, and the latter is equivalent to $\xi\in N(\Sigma_{p})$ (Property (3)). We adapt this argument to the general setting of Alexandrov spaces, with the help of technical results of Perelman [29].
@@
For $a,b\in X$, we denote by $b^{\prime}_{a}\in\Sigma_{a}$ the direction of a shortest path from $a$ to $b$. We also denote by $B(p,r)$ and $\partial B(p,r)$ the open $r$-ball and $r$-sphere around $p$, respectively. Suppose $r$ is small enough and let $x\in\partial B(p,r)$. Then the point $p$ is a $1$-strainer at $x$, that is, there exists a direction at $x$ making an angle almost $\pi$ with any $p^{\prime}_{x}$. Next choose $\varepsilon>0$ small enough compared to $r$. Then the pair of points $(p,x)$ is a $2$-strainer at any $y\in\partial B(p,r)\cap\partial B(x,\varepsilon)$, i.e., $p$ and $x$ are $1$-strainers at $y$ and $p^{\prime}_{y}$ and $x^{\prime}_{y}$ are almost orthogonal. In particular, the distance map $(d(p,\cdot),d(x,\cdot))$ is noncritical at $y$ in the sense of [29, 3.1]. See also [5], [6] for the precise definition and properties of strainers.
$p$$x$$\Sigma$$\partial B(p,r)$$\partial B(x,\varepsilon)$
We will show that $\Sigma_{x}$ is homeomorphic to the topological suspension over
$\Sigma:=\partial B(p,r)\cap\partial B(x,\varepsilon).$
For notational simplicity, suppose that $p^{\prime}_{x}$ is unique (the general case is the same by regarding $p^{\prime}_{x}$ as the set of all directions of shortest paths from $x$ to $p$). Let $\eta\in\Sigma_{x}$ be the unique direction farthest from $p^{\prime}_{x}$ and $\zeta\in\Sigma_{x}$ be the unique direction farthest from $\eta$ (note that $\eta$ and $\zeta$ are at distance almost $\pi$). Then $\zeta$ is $\delta(r)$-close to $p^{\prime}_{x}$, where $\delta(r)\to 0$ as $r\to 0$. By the suspension theorem [29, 4.5] (see also [17, 2.8]), $\Sigma_{x}$ is homeomorphic to the topological suspension $S(\partial B(\zeta,\pi/2))$ over $\partial B(\zeta,\pi/2)\subset\Sigma_{x}$, where $\zeta$ is one of the poles of the suspension (the other pole is a unique direction farthest from $\zeta$). Since $\zeta$ is close enough to $p^{\prime}_{x}$, we see that $\partial B(\zeta,\pi/2)$ is homeomorphic to $\partial B(p^{\prime}_{x},\pi/2)$. More precisely, this follows from the fact that the distance functions $d(\zeta,\cdot)$ and $d(p^{\prime}_{x},\cdot)$ are both regular in the same direction to $\eta$ on the $\pi/5$-neighborhood of $\partial B(\zeta,\pi/2)$. Using this fact, one can construct a noncritical function $\phi$ in the sense of [29, 3.1] of the form
$\phi=\min\{\phi_{1}\circ d(\zeta,\cdot),\phi_{2}\circ d(p^{\prime}_{x},\cdot)\}$
where $\phi_{i}$ ($i=1,2$) are Lipschitz functions chosen so that $\phi$ coincides with $d(\zeta,\cdot)$ around $\partial B(\zeta,\pi/2-\pi/10)$ and with $d(p^{\prime}_{x},\cdot)$ around $\partial B(\zeta,\pi/2+\pi/10)$. Then one can apply the fibration theorem [29, 1.4.1] to $\phi$ to show that $\partial B(\zeta,\pi/2-\pi/10)$ and $\partial B(p^{\prime}_{x},\pi/2+\pi/10)$ are homeomorphic. By the fibration theorem again, these are homeomorphic to $\partial B(\zeta,\pi/2)$ and $\partial B(p^{\prime}_{x},\pi/2)$, respectively. Finally, the stability theorem [29, 4.3] implies that $\Sigma$ is homeomorphic to $\partial B(p^{\prime}_{x},\pi/2)$, provided $\varepsilon$ is small enough. Therefore $\Sigma_{x}$ is homeomorphic to $S(\Sigma)$.
Moreover, as explained in Step 1, any point other than the poles of the suspension is not an extremal point, and thus a manifold point. Therefore, $\Sigma_{x}$ is a topological manifold if and only if $\Sigma$ is simply-connected (Property (3)).
@@
Since $r$ is small enough, it follows from the stability theorem that $\partial B(p,r)$ is homeomorphic to $\Sigma_{p}$. In particular, $\partial B(p,r)$ is a homology manifold with isolated nonmanifold points (Property (2)). Since $\varepsilon$ is small enough, it follows from the fibration theorem (with a gluing argument) that $\partial B(p,r)\cap B(x,\varepsilon)$ is homeomorphic to the open cone $C(\Sigma)$. Thus $\Sigma$ is simply-connected if and only if $x$ is a manifold point of $\partial B(p,r)$ (the “only if” part follows from [8], [4]; see also [26, 6.2] and note that $\dim\partial B(p,r)\geq 4$). In summary,
$\Sigma_{x}$ is a topological sphere (equivalently, a topological manifold)
$\iff$ $\Sigma=\partial B(p,r)\cap\partial B(x,\varepsilon)$ is simply-connected
[…]
Since the rescaled sphere $r^{-1}\partial B(p,r)$ converges to $\Sigma_{p}$ as $r\to 0$ and the stability homeomorphism is a Gromov–Hausdorff approximation, the claim follows. ∎
Claim 4.3 and Property (2) imply that $\Sigma_{p}E$ is an extremal subset of $\Sigma_{p}$. Claim 4.3 also shows that $E$ is closed and has no isolated points. By Lemma 4.2 and the additional condition proved in Step 1, we conclude that $E$ is an extremal subset. This completes the proof. ∎